OpenStudy (anonymous):

What is the equation of a parabola with a vertex at (0, 0) and a focus at (0, 6)

OpenStudy (anonymous):

the equation is x^2 = 4py

OpenStudy (anonymous):

p is te focus in this case is p=6

OpenStudy (anonymous):

x^2=4(6xy)?

OpenStudy (anonymous):

yes it is but let me graph to make sure

OpenStudy (anonymous):

x^2=4(6xy) u must remove the x inside the parentesis x^2 = 24y

OpenStudy (anonymous):

Thank you!^^